what is x+2y=9;(17,y)
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

17+2y=9
2y=-8
y=-4
(17,-4)

Answer by stanbon(75887)

what is x+2y=9;(17,y)
---
Solve for "y" if x = 17.
---
17 + 2y = 9
2y = -8
y = -4
-------------
Cheers,
Stan H.